# MODULI: Unlocking Preference Generalization via Diffusion Models for Offline Multi-Objective Reinforcement Learning

### v1

@@ 4.2 Refining Guidance via Preference-Return Normalization Methods
To perform generation under different scales and physical quantities of objectives, it is essential to establish appropriate guiding conditions, making return normalization a critical component. We first consider directly extending the single-objective normalization scheme to the multi-objective case, named Global Normalization. We normalize each objective’s RTG using min-max style, scaling them to a unified range. Specifically, we calculate the global maximum RTG $\bm{g}^{\text{max}}=\left[g_{1}^{\text{max}},\cdots,g_{n}^{\text{max}}\right]$ for each objective in the offline dataset. The same applies to the minimum RTG $\bm{g}^{\text{min}}$. The conditioned RTG $\bm{g}$ are first normalized to $(\bm{g}-\bm{g}^{\text{min}})/(\bm{g}^{\text{max}}-\bm{g}^{\text{min}})\in[0,1]^{n}$, then concatenated with the preference $\bm{\omega}$ as inputs. During deployment, we aim to generate $\bm{x}(\tau)$ to align the given preference $\bm{\omega}_{\text{target}}$ and maximized scalarized return using the highest RTG condition $\bm{g}=\bm{1}^{n}$. Note that we compute the global extremum values based on the offline dataset with different preferences, which are not necessarily the reachable min/max value for each objective. Also, some objectives are inherently conflicting and can’t maximized at the same time. As a result, direct guidance using the target conditions of $\bm{y}=[\bm{\omega}_{\text{target}},\bm{1}^{n}]$ may be unachievable for some preferences, leading to a performance drop. To address these issues, we match preferences with vector returns and then propose two novel normalization methods that adapt to multi-objective scenarios to achieve preference-related return normalization, providing more refined guidance: Preference Predicted Normalization and Neighborhood Preference Normalization:
@@
To avoid the issue in Global Normalization, we hope to obtain corresponding RTG conditioned on the given preference. The dataset contains a discrete preference space and it is not possible to directly obtain the maximum RTG for each preference. Therefore, we propose Preference Predicted Normalization training a generalized return predictor additionally $\bm{g}^{\text{max}}(\bm{\omega})=\bm{R}_{\psi}(\bm{\omega})$ to capture the maximum expected return achievable for a given preference $\bm{\omega}$. Specifically, we first identify all undominated solution trajectories $\tau_{P}\in D_{P}$, where $D_{P}\subseteq D$ and $P$ demotes Pareto front. Then the generalized return predictor is trained by minimizing the return prediction error conditioned on the preference $\bm{\omega}$:
$\mathcal{L}(\psi)=-\mathbb{E}_{(\bm{\omega},\hat{\bm{g}})\sim D_{P}}\left[\|\bm{R}_{\psi}(\bm{\omega})-\hat{\bm{g}}\|^{2}_{2}\right],$ (7)
@@
Overall, we empirically find that this method can achieve consistently good performance across datasets of various qualities. We leave the comparison of the three normalization methods in the section 5.3 and appendix C.
@@
Now, MODULI can accurately generate desired trajectories with arbitrary preference and RTG in the CFG manner according to eq. 5. Then, we train an additional inverse dynamics model $a_{t}=h(s_{t},s_{t+1})$ to extract the action $a_{t}$ to be executed from generated $\bm{x}$. It is worth noting that we fixed the initial state as $\bm{x}^{k}[0]=\bm{x}^{0}[0]$ to enhance the consistency of the generated results.
@@ 4.3 Enhancing OOD Preference Generalization with Sliding Guidance
After using diffusion models with appropriate normalization, MODULI can exhibit remarkable expressive capabilities, fully covering the ID distribution in the dataset. However, when faced with OOD preferences, it tends to conservatively generate the trajectory closest to the ID preferences. This may indicate that the diffusion model has overfitted the dataset rather than learning the underlying distribution. To enhance generalization to OOD preferences, inspired by [14] and [11], we propose a novel sliding guidance mechanism. After training the diffusion models $p_{\theta}$, We additionally trained a plug-and-play slider adapter with the same network structure of diffusion models, which learns the latent direction of preference changes. This enables controlled continuous concept during generation, actively adjusting the target distribution, thus avoiding simple memorization. We define the adapter as $p_{\theta^{*}}$. When conditioned on $c$, this method boosts the possibility of attribute $c_{+}$ while reduces the possibility of attribute $c_{-}$ according to original model $p_{\theta}$:
$p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)\leftarrow p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta}^{*}\left(c_{+}\mid\bm{x}^{0}(\tau)\right)}{p_{\theta}^{*}\left(c_{-}\mid\bm{x}^{0}(\tau)\right)}\right)^{\eta}$ (11)
Based on eq. 11, we can derive a direct fine-tuning scheme that is equivalent to modifying the noise prediction model. The derivation process is provided in the appendix D:
$\epsilon_{\theta^{*}}\left(\bm{x}^{t},c,t\right)\leftarrow\epsilon_{\theta}^{*}\left(\bm{x}^{t},c,t\right)+\eta\left(\epsilon_{\theta}\left(\bm{x}^{t},c_{+},t\right)-\epsilon_{\theta}\left(\bm{x}^{t},c_{-},t\right)\right)$ (12)
@@
The score function proposed in eq. 12 justifies the distribution of the condition $c$ by converging towards the positive direction $c_{+}$ while diverging from the negative direction $c_{-}$. Therefore, we can let the $\theta^{*}$ model capture the direction of preference changes, that is, we hope that the denoising process of $\epsilon_{\theta}^{*}(x^{t},c,t)$ at each step meets the unit length shift, i.e. $\epsilon_{\theta^{*}}(x^{t},c,t)=\frac{[c_{\theta}(x^{t},c_{+},t)-c_{\theta}(x^{t},c_{-},t)]}{c_{+}-c_{-}}$. In MORL tasks, the condition $c$ is $\omega$, and $c^{+}$ and $c^{-}$ are the changes in preference $\Delta\omega$ in the positive and negative directions, respectively. We minimize the loss $\mathcal{L}(\theta^{*})$ to train the adapter:
@@
$\mathcal{L}(\theta^{*})=\mathbb{E}_{(\bm{\omega},\bm{\Delta\omega})\sim D}\left[\left\|\epsilon_{\theta^{*}}(\bm{x}^{t},\bm{\omega},t)-\frac{[\epsilon_{\theta}(\bm{x}^{t},\bm{\omega}+\Delta{\bm{\omega}},t)-\epsilon_{\theta}(\bm{x}^{t},\bm{\omega}-\Delta\bm{\omega},t)]}{2\Delta\bm{\omega}}\right\|_{2}^{2}\right]$ (13)
@@ 5 Experiments
Table 1: HV and SP performance on Complete datasets. (B: behavioral policy). The best scores are emphasized in bold. Please refer to appendix F for the full results of the Complete datasets at the expert and amateur levels.
@@
We evaluate MODULI on D4MORL [50], an Offline MORL benchmark consisting of offline trajectories from 6 multi-objective MuJoCo environments. D4MORL contains two types of quality datasets: Expert dataset collected by pre-trained expert behavioral policies, and Amateur dataset collected by perturbed behavioral policies. We named the original dataset of D4MORL as the Complete dataset, which is used to verify expressiveness ability to handle multi-objective tasks. Additionally, to evaluate the algorithm’s generalization ability to out-of-distribution (OOD) preferences, similar to fig. 1, we collected two types of preference-deficient datasets namely Shattered and Narrow. We compare MODULI with various strong offline MORL baselines: BC(P), MORvS(P) [12], MODT(P) [6] and MOCQL(P) [25]. These algorithms are modified variants of their single-objective counterparts and well-studied in PEDA [50]. MODT(P) and MORvS(P) view MORL as a preference-conditioned sequence modeling problem with different architectures. BC(P) performs imitation learning through supervised loss, while MOCQL(P) trains a preference-conditioned Q-function based on temporal difference learning. Further details of datasets and baselines are described in appendix A.
@@
We use two commonly used metrics HV and SP for evaluating the empirical Pareto front and an additional metric RD for evaluating the generalization of OOD preference regions: 1 Hypervolume (HV) $\mathcal{H}(P)$ measures the space enclosed by solutions in the Pareto front $P$: $\mathcal{H}(P)=\int_{\mathbb{R}^{n}}{\mathbf{1}_{H(P)}(\bm{z})\mathrm{d}\bm{z}}$, where $H(P)=\{\bm{z}\in\mathbb{R}^{n}|\exists i:1\leq i\leq|P|,\bm{r_{0}}\preceq\bm{z}\preceq P(i)\}$, $\bm{r_{0}}$ is a reference point determined by the environment, $\preceq$ is the dominance relation operator and $\mathbf{1}_{H(P)}$ equals 1 if $\bm{z}\in H(P)$ and 0 otherwise. Higher HV is better, implying empirical Pareto front expansion. 2 Sparsity (SP) $\mathcal{S}(P)$ measures the density of the Pareto front $P$: $\mathcal{S}(P)=\frac{1}{|P|-1}\sum_{i=1}^{n}\sum_{k=1}^{|P|-1}(\tilde{P}_{i}(k)-\tilde{P}_{i}(k+1))^{2}$, where $\tilde{P}_{i}$ represents the sorted list of values of the $i$-th target in $P$ and $\tilde{P}_{i}(k)$ is the $k$-th value in $\tilde{P}_{i}$. Lower SP, indicating a denser approximation of the Pareto front, is better. 3 Return Deviation (RD) $\mathcal{R}(O)$ measures the discrepancy between the solution set $O$ in the out-of-distribution preference region and the maximum predicted value: $\mathcal{R}(O)=\frac{1}{|O|}\sum_{k=1}^{|O|}\|O(k)-\bm{R}_{\psi}(\bm{\omega}_{O(k)})\|_{2}^{2}$, where $\bm{R}_{\psi}$ is pretrained generalized return predictor mentioned in section 4.2, $\bm{\omega}_{O(k)}$ is the corresponding preference of $O(k)$. Lower RD is better, as it indicates that the performance of the solution is closer to the predicted maximum under OOD preference, suggesting better generalization ability. Following the [50], we uniformly select 501 preference points for $2obj$ environments and 325 points for $3obj$ environments. For each experiment, we use 3 random seeds and report the mean and standard deviation.
Table 2: HV, SP and RD on Shattered / Narrow datasets. The best scores are emphasized in bold. Refer to appendix F for full results.
@@
5.1 Evaluating Expressiveness on Complete Datasets
We first compare MODULI with various baselines on the Complete datasets. In this dataset version, the trajectories have comprehensive coverage preferences, and the policy needs sufficient expressive capability to cover complex distributions under a wide range of preferences. As shown in table 1, most of the baselines (MOCQL, MOBC, MODT) exhibit sub-optimal performance, fail to achieve HV close to behavioral policies, and are prone to sub-preference space collapses that do not allow for dense solutions. MORvS can exhibit better HV and SP but fails to achieve performance competitive with behavioral policies due to the poor expressiveness of the context policy. Overall, MODULI achieves the most advanced policy performance according to HV, and the most dense solution with the lowest SP. We visualize the empirical Pareto front in fig. 3, demonstrating that MODULI is a good approximator under different data qualities. In the expert dataset, MODULI accurately replicates the behavioral policies. In the amateur dataset, due to the inclusion of many suboptimal solutions, MODULI significantly expands the Pareto front, demonstrating its ability to stitch and guide the generation of high-quality trajectories.
@@ 5.2 Evaluating Generalization on Shattered and Narrow Datasets
OOD Generalization Visualization In fig. 3, we visualize the actual solutions obtained under OOD preferences. Gray points represent the trajectories in the dataset, where the preference of the dataset is incomplete. We observe that BC(P) exhibits policy collapse when facing OOD preferences, showing extremely low returns and a mismatch with the preferences. RvS(P) demonstrates limited generalization ability, but most solutions still do not meet the preference requirements. Only MODULI can produce correct solutions even when facing OOD preferences. In the Shattered dataset, MODULI can patch partially OOD preference regions, while in the Narrow dataset, MODULI can significantly extend the Pareto front to both sides.
Ablation of Sliding Guidance We conducted an ablation study of the sliding guidance mechanism on 9 tasks within Narrow datasets. The results are presented in the radar chart in fig. 5. We found that in most tasks, MODULI with slider exhibits higher HV and lower RD, indicating leading performance in MO tasks. At the same time, the presence or absence of the slider has little impact on SP performance, indicating that sliding guidance distribution does not impair the ability to express the ID preferences. Besides, we visualized the empirical Pareto front in fig. 5 for Hopper-Narrow dataset. The sliding bootstrap mechanism significantly extends the Pareto set of OOD preference regions, exhibiting better generalization. Due to space limits, we provide more comparative experiments on the different guidance mechanisms (CG, CG+CFG) in appendix G.

### v2

@@ 4.2 Refining Guidance via Preference-Return Normalization Methods
To perform generation under different scales and physical quantities of objectives, it is essential to establish appropriate guiding conditions, making return normalization a critical component. We first consider directly extending the single-objective normalization scheme to the multi-objective case, named Global Normalization. We normalize each objective’s RTG using min-max style, scaling them to a unified range. Specifically, we calculate the global maximum RTG $\bm{g}^{\text{max}}=\left[g_{1}^{\text{max}},\cdots,g_{n}^{\text{max}}\right]$ for each objective in the offline dataset. The same applies to the minimum RTG $\bm{g}^{\text{min}}$. The conditioned RTG $\bm{g}$ are first normalized to $(\bm{g}-\bm{g}^{\text{min}})/(\bm{g}^{\text{max}}-\bm{g}^{\text{min}})\in[0,1]^{n}$, then concatenated with the preference $\bm{\omega}$ as inputs. During deployment, we aim to generate $\bm{x}(\tau)$ to align the given preference $\bm{\omega}_{\text{target}}$ and maximized scalarized return using the highest RTG condition $\bm{g}=\bm{1}^{n}$. Note that we compute the global extremum values based on the offline dataset with different preferences, which are not necessarily the reachable min/max value for each objective. Also, some objectives are inherently conflicting and can’t maximized at the same time. As a result, direct guidance using the target conditions of $\bm{y}=[\bm{\omega}_{\text{target}},\bm{1}^{n}]$ is unachievable for some preferences, leading to a performance drop. To address these issues, we match preferences with vector returns and then propose two novel normalization methods that adapt to multi-objective scenarios to achieve preference-related return normalization, providing more refined guidance: Preference Predicted Normalization and Neighborhood Preference Normalization:
@@
To avoid the issue in Global Normalization, we hope to obtain corresponding RTG conditioned on the given preference. The dataset contains a discrete preference space and it is not possible to directly obtain the maximum RTG for each preference. Therefore, we propose Preference Predicted Normalization training a generalized return predictor additionally $\bm{g}^{\text{max}}(\bm{\omega})=\bm{R}_{\psi}(\bm{\omega})$ to capture the maximum expected return achievable for a given preference $\bm{\omega}$. Specifically, we first identify all undominated solution trajectories $\tau_{P}\in D_{P}$, where $D_{P}\subseteq D$ and $P$ denotes Pareto front. Then the generalized return predictor is trained by minimizing the return prediction error conditioned on the preference $\bm{\omega}$:
$\mathcal{L}(\psi)=\mathbb{E}_{(\bm{\omega},\hat{\bm{g}})\sim D_{P}}\left[\|\bm{R}_{\psi}(\bm{\omega})-\hat{\bm{g}}\|^{2}_{2}\right],$ (9)
@@
Overall, we empirically find that the NPN method achieves consistently good performance across datasets of varying quality, making it the preferred method. On the other hand, PPN, which requires training an NN to estimate the return, performs well only in high-quality datasets. We provide depth analysis of the three normalization methods in the Section 5.3 and Appendix C
@@
Now, MODULI can accurately generate desired trajectories with arbitrary preference and RTG in the CFG manner according to Equation 7. Then, we train an additional inverse dynamics model $a_{t}=h(s_{t},s_{t+1})$ to extract the action $a_{t}$ to be executed from generated $\bm{x}$. It is worth noting that we fixed the initial state as $\bm{x}^{k}[0]=\bm{x}^{0}[0]$ to enhance the consistency of the generated results.
@@ 4.3 Enhancing OOD Preference Generalization with Sliding Guidance
After using diffusion models with appropriate normalization, MODULI can exhibit remarkable expressive capabilities, fully covering the ID distribution in the dataset. However, when faced with OOD preferences, it tends to conservatively generate the trajectory closest to the ID preferences. This may indicate that the diffusion model has overfitted the dataset rather than learning the underlying distribution. To enhance generalization to OOD preferences, inspired by Gandikota et al. (2024) and Du et al. (2020), we propose a novel sliding guidance mechanism. After training the diffusion models $p_{\theta}$, We additionally trained a plug-and-play slider adapter with the same network structure of diffusion models, which learns the latent direction of preference changes. This approach allows for greater control during generation by actively adjusting the target distribution, thereby preventing simple memorization. We define the adapter as $p_{\theta^{*}}$. When conditioned on $c$, this method boosts the possibility of attribute $c_{+}$ while reduces the possibility of attribute $c_{-}$ according to original model $p_{\theta}$:
$p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)\leftarrow p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta}^{*}\left(c_{+}\mid\boldsymbol{x}^{0}(\tau)\right)}{p_{\theta}^{*}\left(c_{-}\mid\boldsymbol{x}^{0}(\tau)\right)}\right)^{\eta}$ (13)
First, we train a diffusion model $p$ to generate trajectories corresponding to ID preferences. Then, we attempt to learn the pattern of change (for example, when the preference shifts from energy efficiency to high speed, the amplitude of the Swimmer’s movements gradually increases). Therefore, our goal is to learn an adapter $p^{*}$, which applies this pattern to OOD preferences to achieve better generalization. In Equation 13, the exponential term represents increasing the likelihood of preference $c+$ and decreasing the likelihood of preference $c-$. $\eta$ represents the guidance strength weight in classifier-free guidance for the diffusion model. Based on Equation 13, we can use an adapter-style method to adjust the original noise prediction model ($p$). Now, the noise prediction combines the noise from the original model and the adapter. Then, we can derive a direct fine-tuning scheme that is equivalent to modifying the noise prediction model. The derivation process is provided in the Appendix D:
$\epsilon_{\theta^{*}}\left(\bm{x}^{t},c,t\right)\leftarrow\epsilon_{\theta}^{*}\left(\bm{x}^{t},c,t\right)+\eta\left(\epsilon_{\theta}\left(\bm{x}^{t},c_{+},t\right)-\epsilon_{\theta}\left(\bm{x}^{t},c_{-},t\right)\right)$
@@
The score function proposed in Equation 14 justifies the distribution of the condition $c$ by converging towards the positive direction $c_{+}$ while diverging from the negative direction $c_{-}$. Therefore, we can let the $\theta^{*}$ model capture the direction of preference changes, that is, we hope that the denoising process of $\epsilon_{\theta}^{*}(x^{t},c,t)$ at each step meets the unit length shift, i.e. $\epsilon_{\theta^{*}}(x^{t},c,t)=\frac{[\epsilon_{\theta}(x^{t},c_{+},t)-\epsilon_{\theta}(x^{t},c_{-},t)]}{c_{+}-c_{-}}$. In MORL tasks, the condition $c$ is $\omega$, and $c^{+}$ and $c^{-}$ are the changes in preference $\Delta\omega$ in the positive and negative directions, respectively. We minimize the loss $\mathcal{L}(\theta^{*})$ to train the adapter:
@@
$\mathcal{L}(\theta^{*})=\mathbb{E}_{(\bm{\omega},\bm{\Delta\omega})\sim D}\left[\left\|\epsilon_{\theta^{*}}(\bm{x}^{t},\bm{\omega},t)-\frac{[\epsilon_{\theta}(\bm{x}^{t},\bm{\omega}+\Delta{\bm{\omega}},t)-\epsilon_{\theta}(\bm{x}^{t},\bm{\omega}-\Delta\bm{\omega},t)]}{2\Delta\bm{\omega}}\right\|_{2}^{2}\right]$
@@ 5 Experiments
We evaluate MODULI on D4MORL (Zhu et al., 2023a), an Offline MORL benchmark consisting of offline trajectories from 6 multi-objective MuJoCo environments. D4MORL contains two types of quality datasets: Expert dataset collected by pre-trained expert behavioral policies, and Amateur dataset collected by perturbed behavioral policies. We named the original dataset of D4MORL as the Complete dataset, which is used to verify expressiveness ability to handle multi-objective tasks. Additionally, to evaluate the algorithm’s generalization ability to out-of-distribution (OOD) preferences, similar to Figure 1, we collected two types of preference-deficient datasets, namely Shattered and Narrow. We refer to the Section A.2 for the dataset generation.
  • Shattered: Simulates incomplete preference distribution by removing part of the trajectories from the Complete dataset, creating regions with missing preferences. It primarily evaluates the interpolation OOD generalization ability, assessing how well the algorithm generalizes when encountering unseen preferences with some missing.
  • Narrow: Similar to the Shattered dataset, but it mainly evaluates the extrapolation OOD generalization ability. The Narrow dataset is generated by removing a portion of the trajectories from both ends of the Complete dataset, resulting in a narrower preference distribution.
We compare MODULI with various strong offline MORL baselines: BC(P), MORvS(P) (Emmons et al., 2021), MODT(P) (Chen et al., 2021) and MOCQL(P) (Kumar et al., 2020). These algorithms are modified variants of their single-objective counterparts and well-studied in PEDA (Zhu et al., 2023a). MODT(P) and MORvS(P) view MORL as a preference-conditioned sequence modeling problem with different architectures. BC(P) performs imitation learning through supervised loss, while MOCQL(P) trains a preference-conditioned Q-function based on temporal difference learning. Further details of datasets and baselines are described in Appendix A.
@@
We use two commonly used metrics HV and SP for evaluating the empirical Pareto front and an additional metric RD for evaluating the generalization of OOD preference regions: 1 Hypervolume (HV) $\mathcal{H}(P)$ measures the space enclosed by solutions in the Pareto front $P$: $\mathcal{H}(P)=\int_{\mathbb{R}^{n}}{\mathbf{1}_{H(P)}(\bm{z})\mathrm{d}\bm{z}}$, where $H(P)=\{\bm{z}\in\mathbb{R}^{n}|\exists i:1\leq i\leq|P|,\bm{r_{0}}\preceq\bm{z}\preceq P(i)\}$, $\bm{r_{0}}$ is a reference point determined by the environment, $\preceq$ is the dominance relation operator and $\mathbf{1}_{H(P)}$ equals 1 if $\bm{z}\in H(P)$ and 0 otherwise. Higher HV is better, implying empirical Pareto front expansion. 2 Sparsity (SP) $\mathcal{S}(P)$ measures the density of the Pareto front $P$: $\mathcal{S}(P)=\frac{1}{|P|-1}\sum_{i=1}^{n}\sum_{k=1}^{|P|-1}(\tilde{P}_{i}(k)-\tilde{P}_{i}(k+1))^{2}$, where $\tilde{P}_{i}$ represents the sorted list of values of the $i$-th target in $P$ and $\tilde{P}_{i}(k)$ is the $k$-th value in $\tilde{P}_{i}$. Lower SP, indicating a denser approximation of the Pareto front, is better. 3 Return Deviation (RD) $\mathcal{R}(O)$ measures the discrepancy between the solution set $O$ in the out-of-distribution preference region and the maximum predicted value: $\mathcal{R}(O)=\frac{1}{|O|}\sum_{k=1}^{|O|}\|O(k)-\bm{R}_{\psi}(\bm{\omega}_{O(k)})\|_{2}^{2}$, where $\bm{R}_{\psi}$ is pretrained generalized return predictor mentioned in Section 4.2, $\bm{\omega}_{O(k)}$ is the corresponding preference of $O(k)$. Lower RD is better, as it indicates that the performance of the solution is closer to the predicted maximum under OOD preference, suggesting better generalization ability. Following the Zhu et al. (2023a), we uniformly select 501 preference points for $2obj$ environments and 325 points for $3obj$ environments. For each experiment, we use 3 random seeds and report the mean and standard deviation.
Table 2: HV, SP and RD on Shattered / Narrow datasets. The best scores are emphasized in bold. See Appendix G for full results.
@@
5.1 Evaluating Expressiveness on Complete Datasets
We first compare MODULI with various baselines on the Complete datasets. In this dataset version, the trajectories have comprehensive coverage preferences, and the policy needs sufficient expressive capability to cover complex distributions under a wide range of preferences. As shown in Table 1, most of the baselines (MOCQL, MOBC, MODT) exhibit sub-optimal performance, fail to achieve HV close to behavioral policies, and are prone to sub-preference space collapses that do not allow for dense solutions. MORvS can exhibit better HV and SP but fails to achieve performance competitive with behavioral policies due to the poor expressiveness of the context policy. Overall, MODULI achieves the most advanced policy performance according to HV, and the most dense solution with the lowest SP. We visualize the empirical Pareto front in Figure 2, demonstrating that MODULI is a good approximator under different data qualities. In the expert dataset, MODULI accurately replicates the behavioral policies. In the amateur dataset, due to the inclusion of many suboptimal solutions, MODULI significantly expands the Pareto front, demonstrating its ability to stitch and guide the generation of high-quality trajectories.
@@ 5.2 Evaluating Generalization on Shattered and Narrow Datasets
OOD Generalization Visualization In Figure 3, we visualize the actual solutions obtained under OOD preferences. Gray points represent the trajectories in the dataset, where the preference of the dataset is incomplete. We observe that BC(P) exhibits policy collapse when facing OOD preferences, showing extremely low returns and a mismatch with the preferences. RvS(P) demonstrates limited generalization ability, but most solutions still do not meet the preference requirements. Only MODULI can produce correct solutions even when facing OOD preferences. In the Shattered dataset, MODULI can patch partially OOD preference regions, while in the Narrow dataset, MODULI can slightly extend the Pareto front to both sides.
Ablation of Sliding Guidance We conducted an ablation study of the sliding guidance mechanism on 9 tasks within Narrow datasets. The results are presented in the radar chart in Figure 4. We found that in most tasks, MODULI with slider exhibits higher HV and lower RD, indicating leading performance in MO tasks. At the same time, the presence or absence of the slider has little impact on SP performance, indicating that sliding guidance distribution does not impair the ability to express the ID preferences. Besides, we visualized the empirical Pareto front in Figure 5 for Hopper-Narrow dataset. The sliding bootstrap mechanism significantly extends the Pareto set of OOD preference regions, exhibiting better generalization. Due to space limits, we provide more comparative experiments on the different guidance mechanisms (CG, CG+CFG) in Appendix H.